# Decentralized Prediction Markets and Sports Books

### v3

@@ 1 Introduction
Decentralized Finance (DeFi) is a novel paradigm utilizing blockchain for intermediating financial transactions. Without utilizing traditional financial intermediaries, DeFi companies provide services in several areas of finance including, but not limited to, lending and borrowing, insurance underwriting, and trading. Herein, we construct a novel formulation for prediction markets within this decentralized paradigm. These markets allow for traders to bet on external outcomes (e.g., sports scores or economic data). Specifically, we formulate these decentralized markets via automated market makers (AMMs). Historically, prediction market AMMs have been considered in a centralized paradigm. However, with the advent of blockchain, decentralized AMMs for digital assets have been created and grown in prominence. These prediction markets function wherein bettors transact against some liquidity pool which takes the other side of every trade. As we will formulate below, and as is commonly formulated for decentralized AMMs, the cost of betting in these markets is found by balancing the liquidity available for all possible outcomes according to mathematical formulas. In contrast to the centralized prediction markets, these decentralized AMMs permit any investor to pool resources into the AMM to assist in making the market; in exchange, these investors share in the profits of the AMM (via fees collected).
The organization of this paper is as follows. The motivation for studying decentralized prediction market AMMs is provided within Section 1.1. A review of the literature on prediction markets and AMMs is provided within Section 1.2. We conclude the introduction with an overview of the primary contributions of this work in Section 1.3. A summary and generalization of prior works on prediction markets is provided within Section 2. This discussion is followed by our proposed structure for a prediction market AMM in an axiomatic framework within Section 3.1. With that discussion, we formulate meaningful mathematical and financial properties for the cost function and pricing oracle in Section 3.2. For instance, we demonstrate that there natively exists bid and ask pricing oracles with a pricing probability measure on the underlying measurable space sandwiched between these prices. In Section 4, we study the problem of decentralized liquidity provision for these prediction markets. In particular, we discuss how to provide liquidity so as to decrease price slippage during trading. In Section 5, we present a discussion of how to explicitly collect fees on bets in this market. Notably, due to the general construction utilized herein, these fees are assessed in a non-trivial manner so as to retain the important financial properties of the AMM. Finally, in Section 6, we consider two numerical case studies. First, we provide an empirical study of the behavior of an AMM when applied to data on sports betting for the 2023 Super Bowl. Second, we apply a prediction market AMMs for the purposes of quoting financial options prices. We summarize and conclude in Section 7.
@@ 1.1 Motivation
DeFi has the opportunity to democratize finance insofar as it opens the positions of financial intermediaries to individual investors. Prior to the “crypto winter” in May 2022, the value of all DeFi projects reached a high of \$180B. [^1] https://tradingview.com/markets/cryptocurrencies/global-charts/ AMMs offer a prominent example of a successful DeFi application; these entities create spot markets between digital assets by holding liquidity pools in an equilibrium. As a decentralized market, these AMMs allow investors to pool their own assets into the market in exchange for a fraction of the fees being collected from trading with the pool.
@@
More specifically, once we define the notion of the prediction market, we want to consider the properties of these AMMs, i.e., for the cost assessed to bettors and the quoted pricing measure. Though many of these properties have been studied previously in the centralized market setting (see the literature review in Section 1.2 below), all works the authors’ are aware of require a finite set of possible outcomes and without any explicit consideration for the impact of liquidity on market behavior. Herein, we generalize such markets to allow for bets to be placed on general probability spaces with the explicit dependence on the cash reserves. Furthermore, due to the possibility of investing in the AMM, it is important to quantify the fees charged to traders. As highlighted in, e.g., [7], naively defining transaction costs can lead to unforeseen consequences to the profits of the AMM. Therefore, to avoid obvious mis-pricing, we also investigate these issues within the setting of prediction markets; prior works avoid these issues due to the centralized nature of the market.
@@ 2 Generalized Structure for Automated Market Makers
Remark 1.
Herein we introduced the AMM and, therefore, the prediction market as a betting market on the real line (see, e.g., [19]). The more typical case for prediction markets incorporates finite probability spaces only. We will frequently utilize the finite probability space $(\Omega_{N},2^{\Omega_{N}},\mathbb{P}_{N})$ with $\mathbb{P}_{N}(\omega)=\frac{1}{N}$ for every $\omega\in\Omega_{N}$ and $\operatorname{card}\Omega_{N}=N\in\mathbb{N}$ in the examples within this work.
The purpose of an automated market maker is to act as a liquidity provider so that individuals can place bets in the prediction market. To place a bet with a payout of $x\in L^{\infty}$, the bettor must make a payment of
$C(x;\pi,L)=\inf\{c\in[\operatorname*{ess\,inf}x,\operatorname*{ess\,sup}x]\;|\;U(\pi+x-\operatorname*{ess\,inf}x,L+c-\operatorname*{ess\,inf}x)\geq U(\pi,L)\}$ (2.1)
@@
The purchasing price $C(x;\pi,L)$ for some bet $x\in L^{\infty}$ can be viewed as a utility indifference price (see e.g. [10]). Specifically, assuming $U$ is strictly increasing in its second (cash reserves) input and sufficiently continuous, we can rewrite $c=C(x;\pi,L)$ as the unique solution to $U(\pi+x-\operatorname*{ess\,inf}x;L+c-\operatorname*{ess\,inf}x)=u(\pi,L)$ provided it exists (i.e., the AMM has sufficient cash reserves to trade the bet $x$). Herein, we choose to follow the optimization formulation of (2.1) as existence and uniqueness of the cost are trivial for all $(x;\pi,L)\in L^{\infty}\times L^{\infty}_{+}\times\mathbb{R}_{+}$.
@@
for any bet $x\in L^{\infty}$, betting position $\pi\in L^{\infty}_{+}$ and cash reserves $L\geq 0$. It is desirable that $P^{b}(\cdot,;\pi,L)\leq P^{a}(\cdot,;\pi,L)$, and that there exists a probability measure $\mathbb{Q}$ such that the bid and ask prices sandwich an expectation with respect to $\mathbb{Q}$.
@@
This also selects the unique measure $\mathbb{Q}$, the expectation with respect to which must then be equal to $P$, i.e., $\mathbb{Q}(A):=P(\mathbb{I}_{A};\pi,L)$ for any $A\in\mathcal{F}$. As such, we can view $\mathbb{Q}$ as a pricing measure, providing a ‘quoted price’ for $x\in L^{\infty}$ via $\mathbb{E}^{\mathbb{Q}}[x]$. Notably, this quoted price differs from $C(x;\pi,L)$ due to the price slippage inherent in the form of $C$ given in (2.1).
@@
$C^{H}(x;\pi,L)=\frac{1}{\gamma}\log\left(\frac{\mathbb{E}[\exp(\gamma[\pi+x])]}{\mathbb{E}[\exp(\gamma\pi)]}\right).$
Recall from Remark 2 that the cost $C^{H}$ can be viewed as a utility indifference price. Specifically writing $U(\pi,L)=u(L\mathbf{1}-\pi)$, and assuming $u$ is strictly increasing, we can rewrite $c=C^{H}(x;\pi,L)$ as the unique solution to $u(L\mathbf{1}+c\mathbf{1}-\pi-x)=u(L\mathbf{1}-\pi)$.
As mentioned within Remark 1, these prediction markets are often considered for finite probability spaces only. Consider now the finite probability space $(\Omega_{N},2^{\Omega_{N}},\mathbb{P}_{N})$ as constructed within that remark. Hanson’s AMM is more commonly written as the integral of the marginal costs [6], i.e., for $x\in\mathbb{R}$
@@ 3 Liquidity-Based Automated Market Makers
As discussed above, within this section we wish to study those automated market makers of the form $U(\pi,L):=u(L\mathbf{1}-\pi)$ for some utility function $u:L^{\infty}_{+}\to\mathbb{R}\cup\{-\infty\}$. [^6] In the subsequent sections we will explicitly provide the domain $\operatorname{dom}u:=\{x\in L^{\infty}_{+}\;|\;u(x)>-\infty\}$ of $u$. Within Section 3.1, we propose some basic axioms that any such liquidity-based automated market maker should satisfy. The implications of these axioms on bets are investigated within Section 3.2. To simplify notation, for the remainder of this work we will let $\Pi:=L\mathbf{1}-\pi$ denote the liquidity remaining for each outcome (contingent on that outcome being realized); we will demonstrate in Section 3.2 that we can always guarantee that $\operatorname*{ess\,inf}\Pi>0$ under the desired axioms (provided that $L_{0}>0$ for the initial cash reserves of the market maker).
@@ 3.1 Construction
    $u$ is upper semi-continuous with respect to the weak* topology; [^8] For any $x$ in the domain of $u$ and any $\epsilon>0$, there exists a neighborhood of $x$ in the weak* topology on $L^{\infty}$ such that for all $y$ in this neighborhood, $u(y)\leq u(x)+\epsilon$.
@@
Note that from Definition 3.1(1)-(2), it follows (see, e.g., [17, Theorem 4.31]) that $u$ is almost surely continuous from above, i.e., if $x_{n}\searrow x$ a.s. then $u(x_{n})\searrow u(x)$.
@@
Consider the finite probability space $(\Omega_{N},2^{\Omega_{N}},\mathbb{P}_{N})$ as in Remark 1 with $\mathbb{P}_{N}(\omega)=\frac{1}{N}$ for every $\omega\in\Omega_{N}:=\{\omega_{1},\dots,\omega_{N}\}$ with cardinality $N$. Consider the logarithmic utility function $u(\Pi)=\mathbb{E}[\log(\Pi)]=\frac{1}{N}\sum_{i=1}^{N}\log(\Pi(\omega_{i}))$ if $\min\Pi>0$ and $u(\Pi)=-\infty$ otherwise. Then, for any $x\in L^{\infty}$, the cost for the bettor
@@
where +:=$\mathcal{F}_{+}:=\{A\in\mathcal{F}\mid\mathbb{P}(A)>0\}$. It is straightforward to verify that this utility function satisfies all the required properties of an LBAMM as per Definition 3.1. Note that, under the finite probability space $(\Omega_{N},2^{\Omega_{N}},\mathbb{P}_{N})$, if $\epsilon\leq\frac{1}{N}$ then $\bar{u}_{\epsilon}\equiv u$. This utility function structure is implemented within Section 6.2 to construct a simple financial derivatives market.
@@ 3.2 Properties
    $u(\Pi-x+C(x;\Pi)\mathbf{1}-y+C(y;\Pi-x+C(x;\Pi))\mathbf{1})=u(\Pi-x+C(x;\Pi)\mathbf{1})=u(\Pi).$
@@
Let $u:L^{\infty}_{*}\to\mathbb{R}$ be an LBAMM with remaining liquidity $\Pi\in L^{\infty}_{*}$. The associated pricing oracles $P^{a},P^{b}:L^{\infty}\to\mathbb{R}$, defined in (3.5), satisfy $P^{a}(x)\geq P^{b}(x)$ for all $x\in L^{\infty}$. Furthermore, there exists a measure $\mathbb{Q}\ll\mathbb{P}$ such that $P^{b}(x)\leq\mathbb{E}^{\mathbb{Q}}[x]\leq P^{a}(x)$ for all $x\in L^{\infty}$.
@@
Since $C$ is convex by Theorem 3.4, it is straightforward to verify that $P^{a},P^{b}:L^{\infty}\to\mathbb{R}$ are well-defined and $\rho(x):=P^{a}(-x)$ is a coherent risk measure. By the dual representation of coherent risk measures (see, e.g., [17, Corollary 4.34]) , there exists a set of probability measures $\mathcal{Q}\neq\emptyset$ which are absolutely continuous with respect to $\mathbb{P}$, i.e., $\mathcal{Q}\subseteq\{\mathbb{Q}\ll\mathbb{P}\}$, such that $P^{a}(x)\geq\sup\bigl{\{}\mathbb{E}^{\mathbb{Q}}[x]\;|\;\mathbb{Q}\in\mathcal{Q}\bigr{\}}$. Moreover, by construction, $P^{b}(x)\leq\inf\bigl{\{}\mathbb{E}^{\mathbb{Q}}[x]\;|\;\mathbb{Q}\in\mathcal{Q}\bigr{\}}$. We thus conclude that $P^{a}(x)\geq\mathbb{E}^{\mathbb{Q}}[x]\geq P^{b}(x)$ for all $x\in L^{\infty}$ and all $\mathbb{Q}\in\mathcal{Q}$. ∎
@@
    Theorem 3.6 guarantees the existence of a pricing measure $\mathbb{Q}\ll\mathbb{P}$ that sits within the bid-ask spread, i.e., $P^{b}(x)\leq\mathbb{E}^{\mathbb{Q}}[x]\leq P^{a}(x)$ for any $x\in L^{\infty}$. We regard any such measure as a pricing measure as it is consistent with the quoted (bid and ask) prices in the market. Furthermore, we wish to note that if $P^{a}(\mathbb{I}_{A})>0$ for all $A\in\mathcal{F}_{+}$ (i.e., $A\in\mathcal{F}$ with $\mathbb{P}(A)>0$), we can guarantee the existence of an equivalent pricing measure $\mathbb{Q}\sim\mathbb{P}$.

### v4

@@ 1.1 Motivation
Decentralized Finance (DeFi) – the novel paradigm utilizing blockchain for financial intermediation – has the opportunity to democratize finance insofar as it opens the positions of financial intermediaries to individual investors. Prior to the “crypto winter” in May 2022, the value of all DeFi projects reached a high of \$180B. [^1] https://tradingview.com/markets/cryptocurrencies/global-charts/ Automated market makers (AMMs) offer a prominent example of a successful DeFi application; these entities create spot markets between digital assets by holding liquidity pools in an equilibrium. As a decentralized market, these AMMs allow investors to pool their own assets into the market in exchange for a fraction of the fees being collected from trading with the pool.
@@
More specifically, once we define the notion of the prediction market, we want to consider the properties of these AMMs, i.e., for the cost assessed to bettors and the quoted pricing measure. Though many of these properties have been studied previously in the centralized market setting (see the literature review in Section 1.2 below), all works the authors’ are aware of require a finite set of possible outcomes and without any explicit consideration for the impact of liquidity on market behavior. Herein, we generalize such markets to allow for bets to be placed on general probability spaces with the explicit dependence on the cash re
Furthermore, due to the possibility of investing in the AMM, it is important to quantify the fees charged to traders. As highlighted in, e.g., [7], naively defining transaction costs can lead to unforeseen consequences to the profits of the AMM. Therefore, to avoid obvious mis-pricing, we also investigate these issues within the setting of prediction markets; prior works avoid these issues due to the centralized nature of the market.
@@ 1.3 Primary contributions
1.4 Organization of this Paper
The organization of this paper is as follows. A summary and generalization of prior works on prediction markets is provided within Section 2. This discussion is followed by our proposed structure for a prediction market AMM in an axiomatic framework within Section 3.1. With that discussion, we formulate meaningful mathematical and financial properties for the cost function and pricing oracle in Section 3.2. For instance, we demonstrate that there natively exists bid and ask pricing oracles with a pricing probability measure on the underlying measurable space sandwiched between these prices. In Section 4, we study the problem of decentralized liquidity provision for these prediction markets. In particular, we discuss how to provide liquidity so as to decrease price slippage during trading. In Section 5, we present a discussion of how to explicitly collect fees on bets in this market. Notably, due to the general construction utilized herein, these fees are assessed in a non-trivial manner so as to retain the important financial properties of the AMM. Finally, in Section 6, we consider two numerical case studies. First, we provide an empirical study of the behavior of an AMM when applied to data on sports betting for the 2023 Super Bowl. Second, we apply a prediction market AMMs for the purposes of quoting financial options prices. We summarize and conclude in Section 7.
@@ 2 Generalized Structure for Automated Market Makers
Fundamentally, an AMM is a liquidity pool – made up of deposits from liquidity providers – against which traders can execute transactions at prices based on mathematical algorithms without requiring a human-in-the-loop. We take the view from DeFi that AMM algorithms are equivalent to utility functions (see, e.g., [37, 7]); the relationship between AMMs and utility functions was previously explored for prediction markets within [12]. The primary purpose of this manuscript is to explain how prediction market AMMs can be constructed so as to explicitly consider the actions, as well as the risks and rewards, of liquidity providers to guarantee there is liquidity available to cover the costs of the winning bets. The AMM begins accepting bets once it has some amount of liquidity and allows bettors to place bets up until some fixed time or event, at which time the winning bets are paid out. [^3] Herein, we do not explicitly consider the evolution of information as would happen, e.g., for in-game sports betting. In Section 4, we consider how to allow for new liquidity provision (or withdrawals of liquidity by liquidity providers) even after the market has already begun accepting bets. However, since the payoffs of the winning bets come from the liquidity pool, liquidity provision is an inherently risky position; because of this, liquidity providers need to be compensated with fees (as disucssed in Section 5). While bettors have a payoff based on the specific bet placed, the liquidity providers share the fees and any liquidity remaining after all bets are paid off.
@@
The classical goal of a prediction market is to be able to give a price to a new incoming bet $x\in L^{\infty}$. This, of course, depends on the current state of the market $(\pi,L)$. As opposed to a standard prediction market, where the cost function of the bet plays the key role, here we utilize the utility indifference approach from DeFi ([37, 7]). In this way, the AMM must not be worse off after a new bet is placed (and the cost of the bet was collected) than it was before this bet was placed. This in turn allows us to define the cost function of any bet. We proceed as follows. That is, to place a bet with a payout of $x\in L^{\infty}$, the bettor must make a payment of
$C(x;\pi,L)=\inf\{c\in[\operatorname*{ess\,inf}x,\operatorname*{ess\,sup}x]\;|\;U(\pi+x-\operatorname*{ess\,inf}x\mathbf{1},L+c-\operatorname*{ess\,inf}x)\geq U(\pi,L)\},$ (2.1)
@@
The purchasing price $C(x;\pi,L)$ for some bet $x\in L^{\infty}$ can be viewed as a utility indifference price (see e.g. [10]). Specifically, assuming $U$ is strictly increasing in its second (cash reserves) input and sufficiently continuous, we can rewrite $c=C(x;\pi,L)$ as the unique solution to $U(\pi+x-\operatorname*{ess\,inf}x\mathbf{1};L+c-\operatorname*{ess\,inf}x)=u(\pi,L)$ provided it exists (i.e., the AMM has sufficient cash reserves to trade the bet $x$). Herein, we choose to follow the optimization formulation of (2.1) as existence and uniqueness of the cost are trivial for all $(x;\pi,L)\in L^{\infty}\times L^{\infty}_{+}\times\mathbb{R}_{+}$.
@@
for any bet $x\in L^{\infty}$, betting position $\pi\in L^{\infty}_{+}$ and cash reserves $L\geq 0$. It is desirable that $P^{b}(\cdot,;\pi,L)\leq P^{a}(\cdot,;\pi,L)$, and that there exists a probability measure $\mathbb{Q}$ such that the bid and ask prices sandwich an expectation with respect to $\mathbb{Q}$. In general, this sandwiched probability measure $\mathbb{Q}$ will depend on the current state $(\pi,L)$.
@@
This also selects the unique measure $\mathbb{Q}$ (for the fixed $(\pi,L)$), the expectation with respect to which must then be equal to the pricing oracle $P$, i.e., $\mathbb{Q}(A):=P(\mathbb{I}_{A};\pi,L)$ for any $A\in\mathcal{F}$. As such, we can view $\mathbb{Q}$ as a pricing measure, providing a ‘quoted price’ for $x\in L^{\infty}$ via $\mathbb{E}^{\mathbb{Q}}[x]$ under market conditions $(\pi,L)$. Notably, this quoted price differs from $C(x;\pi,L)$ due to the price slippage inherent in the form of $C$ given in (2.1).
@@
$C^{H}(x;\pi,L)=\frac{1}{\gamma}\log\left(\frac{\mathbb{E}[\exp(\gamma(\pi+x))]}{\mathbb{E}[\exp(\gamma\pi)]}\right).$
In this case $\frac{d\mathbb{Q}}{d\mathbb{P}}=\frac{\operatorname{e}^{\gamma\pi}}{\mathbb{E}[\operatorname{e}^{\gamma\pi}]}$ is the pricing measure under market conditions $(\pi,L)$; notably, while $\mathbb{Q}$ does not depend on the existing reserves $L$, it does depend on the existing bets $\pi$. Recall from Remark 1 that the cost $C^{H}$ can be viewed as a utility indifference price. Specifically writing $U(\pi,L)=u(L\mathbf{1}-\pi)$, and assuming $u$ is strictly increasing, we can rewrite $c=C^{H}(x;\pi,L)$ as the unique solution to $u(L\mathbf{1}+c\mathbf{1}-\pi-x)=u(L\mathbf{1}-\pi)$.
Often, prediction markets are only considered with finite probability spaces [23]. Here, and in other examples below, we will utilize the finite probability space $(\Omega_{N},2^{\Omega_{N}},\mathbb{P}_{N})$ with $\mathbb{P}_{N}(\omega)=\frac{1}{N}$ for every $\omega\in\Omega_{N}$ and $\operatorname{card}(\Omega_{N})=N$, for some $N\in\mathbb{N}$. Hanson’s AMM is more commonly written as the integral of the marginal costs [6], i.e., for $x\in\mathbb{R}$,
@@ 3 Liquidity-Based Automated Market Makers
As discussed above, within this section we wish to study those automated market makers of the form $U(\pi,L):=u(L\mathbf{1}-\pi)$ for some utility function $u:L^{\infty}_{+}\to\mathbb{R}\cup\{-\infty\}$. [^7] In the subsequent sections we will explicitly provide the domain $\operatorname{dom}u:=\{x\in L^{\infty}_{+}\;|\;u(x)>-\infty\}$ of $u$. As introduced in Section 2, we again define a prediction market AMM by its utility function $u$ rather than a cost function of a bet $C$; we then derive the cost function using the utility indifference argument. Within Section 3.1, we propose some basic axioms that any such liquidity-based automated market maker should satisfy. The implications of these axioms on bets are investigated within Section 3.2. To simplify notation, for the remainder of this work we will let $\Pi:=L\mathbf{1}-\pi$ denote the liquidity remaining for each outcome (contingent on that outcome being realized); we will demonstrate in Section 3.2 that we can always guarantee that $\operatorname*{ess\,inf}\Pi>0$ under the desired axioms (provided that $L_{0}>0$ for the initial cash reserves of the market maker).
@@ 3.1 Construction
    $u$ is upper semi-continuous on $L^{\infty}_{+}$ with respect to the weak* topology; [^9] For any $x$ in the domain of $u$ and any $\epsilon>0$, there exists a neighborhood of $x$ in the weak* topology on $L^{\infty}$ such that for all $y$ in this neighborhood, $u(y)\leq u(x)+\epsilon$.
@@
Note that from Definition 3.1(1)-(2), it follows (see, e.g., [18, Theorem 4.31]) that $u$ is almost surely continuous from above, i.e., if $x_{n}\searrow x$ a.s. then $u(x_{n})\searrow u(x)$.
@@
Consider the finite probability space $(\Omega_{N},2^{\Omega_{N}},\mathbb{P}_{N})$ as in Example 2.2 with $\mathbb{P}_{N}(\omega)=\frac{1}{N}$ for every $\omega\in\Omega_{N}:=\{\omega_{1},\dots,\omega_{N}\}$ with cardinality $N$. Consider the logarithmic utility function $u(\Pi)=\mathbb{E}[\log(\Pi)]=\frac{1}{N}\sum_{i=1}^{N}\log(\Pi(\omega_{i}))$ if $\min\Pi>0$ and $u(\Pi)=-\infty$ otherwise. As discussed in, e.g., [7], the constant product market maker of Uniswap V2 is equivalent to a logarithmic utility indifference pricing; for this reason we consider this LBAMM to be the prediction market version of Uniswap V2. Then, for any $x\in L^{\infty}$, the cost for the bettor
@@
where +:=$\mathcal{F}_{+}:=\{A\in\mathcal{F}\mid\mathbb{P}(A)>0\}$. It is straightforward to verify that this utility function satisfies all the required properties of an LBAMM as per Definition 3.1. Note that, under the finite probability space $(\Omega_{N},2^{\Omega_{N}},\mathbb{P}_{N})$, if $\epsilon\leq\frac{1}{N}$ then $\bar{u}_{\epsilon}\equiv u$. This utility function structure is implemented within Section 6.2 to construct a simple financial derivatives market. [^12] The gas fees for using a general utility function $\bar{u}_{\epsilon}$ can be controlled by utilizing modern blockchains such as Avalanche and Skale which charge developers a monthly fee for gas-less transactions.
@@ 3.2 Properties
    $u(\Pi-x+C(x;\Pi)\mathbf{1}-y+C(y;\Pi-x+C(x;\Pi)\mathbf{1})\mathbf{1})=u(\Pi-x+C(x;\Pi)\mathbf{1})=u(\Pi).$
@@
Let $u:L^{\infty}_{*}\to\mathbb{R}$ be an LBAMM with remaining liquidity $\Pi\in L^{\infty}_{*}$. The associated pricing oracles $P^{a},P^{b}:L^{\infty}\to\mathbb{R}$, defined in (3.5), satisfy $P^{a}(x)\geq P^{b}(x)$ for all $x\in L^{\infty}$. Furthermore, there exists a measure $\mathbb{Q}\sim\mathbb{P}$ such that $P^{b}(x)\leq\mathbb{E}^{\mathbb{Q}}[x]\leq P^{a}(x)$ for all $x\in L^{\infty}$.
@@
Since $C$ is convex by Theorem 3.4, it is straightforward to verify that $P^{a},P^{b}:L^{\infty}\to\mathbb{R}$ are well-defined and $\rho(x):=P^{a}(-x)$ is a lower semicontinuous coherent risk measure. In fact, following the nomenclature of [17], $\rho$ is also relevant as $P^{a}(\mathbb{I}_{A})>0$ for every $A\in\mathcal{F}_{+}$ (i.e., $A\in\mathcal{F}$ with $\mathbb{P}(A)>0$). [^13] Trivially $P^{a}(\mathbb{I}_{A})=1-P^{b}(\mathbb{I}_{A^{c}})>0$ for any $A\in\mathcal{F}_{+}$ as $P^{b}(\mathbb{I}_{A^{c}})\leq P^{a}(\mathbb{I}_{A^{c}})\leq C(\mathbb{I}_{A^{c}})<1$ by no arbitrage and convexity (Theorem 3.4). By the dual representation of coherent risk measures (see, e.g., [18, Corollary 4.34]) , there exists a set of probability measures $\mathcal{Q}\neq\emptyset$ which are equivalent to $\mathbb{P}$, i.e., $\mathcal{Q}\subseteq\{\mathbb{Q}\sim\mathbb{P}\}$, such that $P^{a}(x)=\sup\bigl{\{}\mathbb{E}^{\mathbb{Q}}[x]\;|\;\mathbb{Q}\in\mathcal{Q}\bigr{\}}$. Moreover, by construction, $P^{b}(x)=\inf\bigl{\{}\mathbb{E}^{\mathbb{Q}}[x]\;|\;\mathbb{Q}\in\mathcal{Q}\bigr{\}}$. We thus conclude that $P^{a}(x)\geq\mathbb{E}^{\mathbb{Q}}[x]\geq P^{b}(x)$ for all $x\in L^{\infty}$ and all $\mathbb{Q}\in\mathcal{Q}$. ∎
@@
    Theorem 3.6 guarantees the existence of an equivalent pricing measure $\mathbb{Q}\sim\mathbb{P}$ that sits within the bid-ask spread, i.e., $P^{b}(x)\leq\mathbb{E}^{\mathbb{Q}}[x]\leq P^{a}(x)$ for any $x\in L^{\infty}$. We remind the reader that this equivalent pricing measure $\mathbb{Q}$ implicitly depends on the remaining liquidity $\Pi\in L^{\infty}_{*}$. We regard any such measure as a pricing measure as it is consistent with the quoted (bid and ask) prices in the market.
@@
We conclude our discussion of the properties of the LBAMM by considering the outcome of a bet from an infinitely liquid and risk-neutral bettor with beliefs $\mathbb{Q}\ll\mathbb{P}$, i.e., who is solving
$\sup_{x\in L^{\infty}}\mathbb{E}^{\mathbb{Q}}[x-C(x;\Pi)\mathbf{1}].$ (3.6)
Such a result serves as a converse to Theorem 3.6 as this bettor has a (finite) optimal bet $x^{*}\in L^{\infty}$ if, and only if, $P^{b}(x;\Pi-x^{*}+C(x^{*};\Pi)\mathbf{1})\leq\mathbb{E}^{\mathbb{Q}}[x]\leq P^{a}(x;\Pi-x^{*}+C(x^{*};\Pi)\mathbf{1})$ for any $x\in L^{\infty}$ so that $\mathbb{Q}$ is sandwiched by $P^{b},P^{a}$. Fundamentally, this sandwich property implies an LBAMM is able to extract the subjective beliefs (i.e., $\mathbb{Q}$) of any risk-neutral bettor. Such a property was first considered within [23] in which the Hanson’s logarithmic scoring rule in a finite probability space was proven to “extract the information implicit in the trades others make with it, in order to infer new rational prices.”
Theorem 3.7.
Let $u:L^{\infty}_{*}\to\mathbb{R}$ be an LBAMM with remaining liquidity $\Pi\in L^{\infty}_{*}$. Consider a risk-neutral bettor with subjective measure $\mathbb{Q}\ll\mathbb{P}$, optimizing (3.6). For any $x^{*}\in\operatorname*{arg\,max}_{x\in L^{\infty}}\mathbb{E}^{\mathbb{Q}}[x-C(x;\Pi)\mathbf{1}]$, it holds that
$P^{b}(x;\Pi-x^{*}+C(x^{*};\Pi)\mathbf{1})\leq\mathbb{E}^{\mathbb{Q}}[x]\leq P^{a}(x;\Pi-x^{*}+C(x^{*};\Pi)\mathbf{1}).$
Furthermore, if $Q\in\mathcal{M}_{*}:=\{\mathbb{Q}\ll\mathbb{P}\;|\;\operatorname*{ess\,inf}\frac{d\mathbb{Q}}{d\mathbb{P}}>0\}$ then $\operatorname*{arg\,max}_{x\in L^{\infty}}\mathbb{E}^{\mathbb{Q}}[x-C(x;\Pi)\mathbf{1}]\neq\emptyset$.
Proof.
By definition of the pricing oracles $P^{a},P^{b}$ as directional derivatives, the sandwich property holds if $\frac{d\mathbb{Q}}{d\mathbb{P}}\in\partial C(\mathbf{0};\Pi-x^{*}+C(x^{*};\Pi)\mathbf{1})$ for $x^{*}\in\operatorname*{arg\,max}_{x\in L^{\infty}}\mathbb{E}^{\mathbb{Q}}[x-C(x;\Pi)\mathbf{1}]$ where $\partial C$ denotes the subdifferential of $C$. By construction of $x^{*}$ as a maximizer, $\mathbb{E}^{\mathbb{Q}}[x^{*}-C(x^{*};\Pi)\mathbf{1}]\geq\mathbb{E}^{\mathbb{Q}}[x-C(x;\Pi)\mathbf{1}]$ for any $x\in L^{\infty}$. Rearranging terms, we recover $C(x;\Pi)-C(x^{*};\Pi)\geq\mathbb{E}^{\mathbb{Q}}[x-x^{*}]$ for any $x\in L^{\infty}$, i.e., $\frac{d\mathbb{Q}}{d\mathbb{P}}\in\partial C(x^{*};\Pi)$. Utilizing path independence (Theorem 3.4(4)), it follows that $\mathbf{0}\in\operatorname*{arg\,max}_{x\in L^{\infty}}\mathbb{E}^{\mathbb{Q}}[x-C(x;\Pi-x^{*}+C(x^{*};\Pi)\mathbf{1})\mathbf{1}]$, i.e., $\frac{d\mathbb{Q}}{d\mathbb{P}}\in\partial C(\mathbf{0};\Pi-x^{*}+C(x^{*};\Pi)\mathbf{1})$.
It remains to show that $\operatorname*{arg\,max}_{x\in L^{\infty}}\mathbb{E}^{\mathbb{Q}}[x-C(x;\Pi)\mathbf{1}]\neq\emptyset$ for $\mathbb{Q}\in\mathcal{M}_{*}$. Let $\mathcal{A}:=\{z\in L^{\infty}\;|\;u(z)\geq u(\Pi)\}\subseteq L^{\infty}_{+}$ be the superlevel set at $u(\Pi)$. By quasiconcavity and upper-semicontinuity, $\mathcal{A}$ is a weak* closed and convex set. Fix $y\in L^{1}_{+}$, then
$\sup_{x\in L^{\infty}}\mathbb{E}[y(x-C(x;\Pi)\mathbf{1})]$ $=\sup_{x\in L^{\infty}}\{\mathbb{E}[yx]\;|\;u(\Pi-x)=u(\Pi)\}=\sup_{x\in L^{\infty}}\{\mathbb{E}[yx]\;|\;u(\Pi-x)\geq u(\Pi)\}$
$=\sup_{x\in L^{\infty}}\{\mathbb{E}[yx]\;|\;\Pi-x\in\mathcal{A}\}=\mathbb{E}[y\Pi]-\inf_{x\in\mathcal{A}}\mathbb{E}[yx].$
By $\mathcal{A}\subseteq L^{\infty}_{+}$, it follows that $\inf_{x\in\mathcal{A}}\mathbb{E}[yx]\geq\inf_{x\in L^{\infty}_{+}}\mathbb{E}[yx]=0$. As a direct consequence, we also have that $\sup_{x\in L^{\infty}}\mathbb{E}[y(x-C(x;\Pi)\mathbf{1})]<\infty$ for any $y\in L^{1}_{+}$. For ease of notation, define :=$\mathcal{B}:=\{x\in L^{\infty}\;|\;\Pi-x\in\mathcal{A}\}$ is weak* closed and convex. Further, we can define the associated indicator function $\delta(x):=\begin{cases}0&\text{if }x\in\mathcal{B}\\ \infty&\text{if }x\not\in\mathcal{B}\end{cases}$ (proper, weak* lower semicontinuous, and convex) and support function $\delta^{*}(y):=\sup_{x\in\mathcal{B}}\mathbb{E}[yx]$ (proper, weak* lower semicontinuous, and convex) for $y\in L^{1}$. By the prior finiteness of the supremum result, $\operatorname{dom}\delta^{*}\supseteq L^{1}_{+}$ and, in particular, $\operatorname{int}\operatorname{dom}\delta^{*}\supseteq L^{1}_{*}:=\{y\in L^{1}_{+}\;|\;\operatorname*{ess\,inf}y>0\}$. Therefore $\partial\delta^{*}(\frac{d\mathbb{Q}}{d\mathbb{P}})\neq\emptyset$ since $\frac{d\mathbb{Q}}{d\mathbb{P}}\in\operatorname{int}\operatorname{dom}\delta^{*}$. By the Fenchel-Young inequality, any subdifferential $x^{*}\in\partial\delta^{*}(\frac{d\mathbb{Q}}{d\mathbb{P}})$ is a maximizer of $\sup_{x\in L^{\infty}}\mathbb{E}^{\mathbb{Q}}[x-C(x;\Pi)\mathbf{1}]$ and the result follows. ∎